OpenStudy (anonymous):

There are 52 cards in a deck. There are 4 eights and 4 nines. What is the probability that if you cut the deck, you'll get an eight or a nine? Express your answer as a decimal number rounded to the hundredths place.

OpenStudy (anonymous):

8/52 = 15.38 % chance

OpenStudy (anonymous):

all you have to do is divide the amount of cards that will satisfy your restriction by the total number of cards i.e 8/52 =2/13